OpenStudy (anonymous):

y=1/3x-4 y=7/3x+4 solving system of equation

OpenStudy (anonymous):

start with \[\frac{1}{3}x-4=\frac{7}{3}x+4\] then multiply by 3 and get \[x-12=7x+12\]

OpenStudy (anonymous):

subtract \(x\) to get \[-12=6x+12\] then subtract \(12\) giviing \[-24=6x\] and finally solve for \(x\)